A geologist finds that a Moon rock whose mass is 9.28 kg has an apparent mass of 6.23 kg when submerged in water. What is the density of the rock?
Answer #1

here,

mass of moon rock , m = 9.28 kg

apparent mass , m' = 6.23 kg

let the density of rock be p

buoyant force = g * ( m - m')

density of water * Volume of mass * g = g * (m - m')

pw * ( m /p) = ( m - m')

1000 * ( 9.28 /p) = ( 9.28 - 6.23)

solving for p

p = 3042.6 Kg/m^3

the density of the rock is 3042.6 kg/m^3